OpenStudy (loser66):

Let G be a group and let K and N be normal subgroups of G. Prove that there is an injective group homomorphism from \(G/K\cap N\) to (G/K) x (G/N) I have homomorphism part done but injective. Please, help

OpenStudy (loser66):

@Alchemista This time, I carefully type the problem. :)

OpenStudy (anonymous):

ya

OpenStudy (loser66):

someone tells me please, how to type x in latex such that it does not appear as \(x\)?

OpenStudy (loser66):

@zepdrix

OpenStudy (anonymous):

cool

OpenStudy (loser66):

don't mess up my post , please,

OpenStudy (loser66):

I defined: \[f: G/K\cap N \rightarrow (G/K)\times (G/N)\\~~~~g~(K\cap N)\rightarrow (gK, gN)\]

OpenStudy (loser66):

Let a, b \(\in G\), then \(a(K\cap N),~~b(K\cap N)\in K\cap N \) and \(a(K\cap N) b(K\cap N)=ab(K\cap N)\)

OpenStudy (mathmath333):

\(\huge\tt \begin{align} \color{black}{x }\end{align}\)

OpenStudy (mathmath333):

\(\huge\tt \begin{align} \color{black}{\times }\end{align}\)

OpenStudy (loser66):

so that we have \[f(ab (K\cap N) =(ab K, ab N)\] but ab K = aK bK and ab N = aN bN so that it is \((aK, aN)(bK, bN)= f(a (K\cap N)*f(b(K\cap N)\) that shows f is homomorphism

OpenStudy (anonymous):

Show that the kernel is trivial.

OpenStudy (loser66):

show me, please

OpenStudy (anonymous):

Show that the only element that gets sent to \((eK, eN)\) is the identity.

OpenStudy (loser66):

Is it not that we have to show f is one - to -one?

OpenStudy (anonymous):

These are equivalent. Another way would be to show that $$ f(a) =f(b) \implies a = b $$

OpenStudy (anonymous):

\(\begin{eqnarray*} f(a) &=& f(b) \\ (aK, aN) &=& (bK, bN) \end{eqnarray*}\) Does this mean that \(a=b\)?

OpenStudy (loser66):

yes,

OpenStudy (anonymous):

Make sure you clarify that this is true since they are equivalence class representatives.